OpenStudy (anonymous):

v=u^2+5as u=6 a=2.5 s=9

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

plug in u first. \[\large u^2 = 6*6\] \[\large 5as = 5*2.5*9\]

OpenStudy (bibby):

Should I expand further?

OpenStudy (anonymous):

148.5 right?

OpenStudy (bibby):

36+45*2.5 = 36+112.5 =148.5 Good job :D

OpenStudy (anonymous):

thnx